OpenStudy (anonymous):

Evaluate ∫ 100 ^e-0.01x dx

OpenStudy (anonymous):

let u = .01x, du = .01dx.

OpenStudy (anonymous):

For clarification, is this your question? \[\int\limits_{}^{}100^{(e-0.01x)}dx\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okie dokie, first step would be to use the rules of exponents to separate the integral into \[\int\limits_{}^{}100^{(e-0.01x)}dx = \int\limits_{}^{}(100^e \times100^{-0.01x})dx=100^e \int\limits_{}^{}100^{-0.01x}dx\] Try solving it from here. Hint: rewrite 100 as a power of 10 and -0.01 as a fraction and see what ya get.